OpenStudy (anonymous):

Factor the trinomial completely: 16p^4-40p^3+25p^2 ^ means exponent. (as you can tell factoring trinomials are not my strong point.)

OpenStudy (anonymous):

16p^4-40p^3+25p^2 =p^2(16p^2-40p+25) =p^2(4p-5)^2